# Data Compression for Fast Online Stochastic Optimization

### v1

@@ 3.4 Performance guarantees
4 Online performance analysis
So far, we have obtained guarantees with respect to the out-of-sample performance, and to the nominal DRO performance with the same information, i.e., both having observed $\mathcal{D}_{t-1}$. However, in online settings, we are interested in robustifying our solution with respect to the subsequent realization of the uncertainty, i.e., we want to compare with an oracle that has observed $\mathcal{D}_{t}$. This notion of regret is common in online learning [3, 44, 9], where the online solution is compared against the best solution in hindsight, and measures how well an algorithm can adapt to anticipated data.
In this section, we thus compute the dynamic regret bound, where we compare our online solution with information up to time $t-1$, against the nominal DRO solution with information up to time $t$. In particular, recall that we solve for the online solution at time $t$ before the new data-point is observed, with the ambiguity set $\mathcal{P}^{K}_{t-1}$. The dynamic regret, defined in equation (9), then calculates the difference between the worst-case performance of our online solution $x^{K}_{t}$, hedged against the updated non-compressed ambiguity set $\mathcal{P}_{t}$, and the optimal DRO value over the same set. Note that the performance of the online solution over the set $\mathcal{P}_{t}$, which is built around the full empirical distribution at time $t$, will differ from its performance over the set $\mathcal{P}^{K}_{t-1}$, which it was optimized against. We present the following results; all proofs are delayed to Appendix A.2.
Lemma 4.1.
Suppose assumptions 3.3 and 3.4 hold. The dynamic regret satisfies
$R(T,K)\leq\frac{1}{T}\sum_{t=1}^{T}\left(\underline{\psi}_{t+1}^{K}+\bar{\psi}_{t}^{K}\right)$ (15)
$\quad+\frac{1}{T}\left(\max_{j\leq J}M_{j}\left(\varepsilon_{0}+\varepsilon_{T}+W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})+\sum_{t=1}^{T}(2\varepsilon_{t-1}+W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1}))\right)\right),$
with $\underline{\psi}_{t}^{K}$ and $\bar{\psi}^{K}_{t}$ as defined in Theorem 3.1.
Note that there are two sources of discrepancy: one from clustering, and one from the difference in time step. The difference in time step results in an offset in the index of $\underline{\psi}_{t+1}^{K}$, as well as the addition of the second set of terms. However, these differences once again boil down to the discrepancy between the distributions involved, for which we have convergence results. We thus obtain the following bounds.
Theorem 4.1 (Dynamic regret bound).
Suppose Assumption 3.3 and the first clause of Assumption 3.4 holds. For finite $T\gg\tau$, with probability $1-\zeta$, the dynamic regret is bounded as
$R(T,K)\leq O\left(\left(\log(\beta_{T}^{-1})/T\right)^{1/d}\right)+O\left(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right),$
where $\beta_{T}$ decays at least sublinearly in $T$, and $\zeta$ scales with $\beta_{T}$.
Corollary 4.1.1.
If we choose $\beta_{t}=O(\exp(-\sqrt{n_{t}}))$ in Assumption 3.3, the dynamic regret is bounded as
$R(T,K)\leq O\left(T^{-\frac{1}{2d}}\right)+O\left(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right),$
with some probability $1-\zeta\geq 1-\sum_{t=0}^{T}\beta_{t}$.
In Theorem 4.1, we assume the Lipschitz condition in order to obtain a clear convergence rate; the regret converges sublinearly to a value at most $O(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star}))$, a factor of the Wasserstein-1 distance between the true distribution and the converged compressed distribution. This is in line with our expectations. Recall that over time, we gain confidence in the empirical measures as approximations of the true distributions. Therefore, both the ambiguity set radii we select and the theoretical distances between the empirical and true distributions converge sublinearly to 0. What remains is then a function of the discrepancy between the converged (true) distributions.
Theorem 4.1 thus consolidates theorems 3.53.6 and 3.7, and shows that the difference arising from the one-time-step information discrepancy converges to 0. Overall, only the choice of $K$ and the clustering algorithm directly affect the asymptotic performance of the online algorithm: they fully control the trade-off between efficiency and optimality. Below, we note this asymptotic behavior, which hold due to ${\mathbf{P}}^{\infty}$-almost sure convergence.
Corollary 4.1.2.
Suppose assumptions 3.3 and 3.4 hold. With probability 1, asymptotically we observe $\lim_{T\to\infty}R(T,K)\leq O(\underline{\psi}_{\star}^{K}+\bar{\psi}_{\star}^{K}),$ and additionally as $K\to\infty$$\lim_{K\to\infty}\lim_{T\to\infty}R(T,K)=0.$
@@ 5 Special case: bounded support
In Section 3, we gave performance guarantees for light-tailed supports $S$, and set the ambiguity set radius $\varepsilon^{K}_{t}$ with the nominal DRO radius $\varepsilon_{t}$, calculated using Theorem 3.1. In this section, for the case of bounded support, we give an explicit formulation of the radius in terms of the diameter of the support $S$. We also derive performance guarantees using a different theoretical approach from that of Section 3: using measure concentration results, we instead increase the ambiguity set radius by a finite value. In this case, the online solution is guaranteed to be an upper bound on the DRO solution, regardless of the curvature of $f$.
@@ 5.1 Wasserstein radius with bounded support
This follows from [11, Proposition A.2] and exploits the fact that bounded distributions are sub-Gaussians. Then, by bounding the expected Wasserstein distance using a function of $d,p,\tilde{\rho}$, and $N$, we obtain the following explicit characterization of the nominal ambiguity radius. For details, see the proofs in the online version of the referenced paper.
@@
For any time step $t$ and $\beta_{t}$, the radius $\varepsilon_{t}$ calculated from Theorem 5.1 can be used to obtain the performance guarantees in Section 3, since distributions with bounded support are trivially light-tailed.
@@ 5.2 Performance guarantees with bounded support
Let $\hat{{\mathbf{P}}}^{K}_{t}$ be the clustered empirical distribution at time $t$, with cluster diameters at most $2\eta_{K}$. For all $t$, we have $d_{W_{p}}(K,t)\leq 2\eta_{K}.$ With probability $1-\beta_{t}$, we have $W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})\leq\varepsilon_{t}+2\eta_{K},$ where $\beta_{t}$ and $\varepsilon_{t}$ are computed as in Theorem 5.1.
@@
Under the assumptions of Theorem 5.1, let the adaptive ambiguity set $\mathcal{P}^{K}_{t-1}$ be defined with $\varepsilon^{K}_{t-1}=\varepsilon_{t-1}+2\eta_{K}$. Then, the solution $x^{K}_{t}$ and optimal value $H^{K}_{t}$ of the compressed DRO problem (6) implies the finite sample performance guarantee
${\mathbf{P}}^{n_{t-1}}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x^{K}_{t})]\leq H^{K}_{t}\right)\geq 1-\beta_{t-1}.$
@@
This follows from Lemma 5.1 and the same logic as nominal DRO results [40, Theorem 3.5]. ∎
In addition, we can derive the following asymptotic result by applying a similar approach as [40, Lemma 3.7].
@@
This lemma dictates that asymptotically, the true distribution lies in a ball around the clustered distribution with radius $4\eta_{K}$; the increased radius is the price to pay due to the limited computational budget available to solve (4), and is directly related to the quality of our approximation $\hat{{\mathbf{P}}}^{K}_{t}$. Clearly, if $K$ is high, then $\eta_{K}$ will be low and our asymptotic estimate will be more accurate; viceversa for low $K$. Following [40, Theorem 3.6], we now derive the following result.
@@
Assume for all $t$, $\varepsilon_{t}$ is computed as in Theorem 5.1, with a sequence $\beta_{t}$ such that $\sum_{t=0}^{\infty}\beta_{t}<\infty$ and $\lim_{t\to\infty}\varepsilon_{t}=0$. Also assume the adaptive ambiguity set $\mathcal{P}^{K}_{t}$ is defined with $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$, and that for all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ satisfy the $M$-Lipschitz condition given in Definition 2.1 with constants $M_{j}$. Then, the sequence of solutions and optimal values $x^{K}_{t}$ and $H^{K}_{t}$ of the compressed DRO problem (6) satisfies
@@ 6 Online algorithms for data-compression
In Section 3.2, we gave requirements for the clustering algorithm. In this section, we provide three online clustering algorithms satisfying the given assumptions. These algorithms are expected to maintain a reasonably accurate approximation of the optimal discretization while minimizing computational effort. Below, we summarize their time and memory complexities.
Table 1: Time and memory complexities of our clustering algorithms. Recall that $n_{t}$ is the total number of data-points at time $t$, $K$ is the number of clusters, and $d$ is the dimension of the data. We also let $I$ be the number of iterations required for the clustering algorithm’s convergence. The values $G$ and $Q$ are defined in the respective algorithms.
@@ 6.1 A bounded coverage of the support (SupCover)
3:assign data-points to clusters $\{C^{k}_{0}\}_{k=1}^{K}$, and compute $\{n^{k}_{0}\}_{k=1}^{K}$, $\{\theta^{k}_{0}\}_{k=1}^{K}$, $\{\bar{u}_{0}^{k}\}_{k=1}^{K}$
@@
6:     $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\}$triangleright assign $\hat{u}$ to cluster $k^{\prime}$ following  (16)
@@
8:     $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$triangleright update centroid
@@
where $\hat{u}$ is the data-point at time $t$. Additionally at each time $t$, the centroid of the cluster that gained a data-point is updated. Notably, we do not need to store $\mathcal{D}_{t}$. By keeping track of the centers, centroids and weights of each cluster, together with a count of the total number of data-points seen, we can perform the above updates with only the newest data-point. After performing the update, this data-point can also be discarded. The algorithm is thus memory efficient; it only requires $O(Kd)$ space.
@@ 6.2 Approximate $k$-centers clustering with warm starts (Reclustering)
2:$A_{0}\leftarrow$ find $K$ centers using $k$-centers on $\mathcal{D}_{0}$
3:assign data-points to clusters $\{C^{k}_{0}\}_{k=1}^{K}$, and compute $\{n^{k}_{0}\}_{k=1}^{K}$, $\{\theta^{k}_{0}\}_{k=1}^{K}$, $\{\bar{u}_{0}^{k}\}_{k=1}^{K}$
@@
8:         $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\}$triangleright assign $\hat{u}$ to cluster $k^{\prime}$ following  (16)
9:         $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$triangleright update centroid
@@
In the complexity analysis, we assume the $k$-centers approximation to be $k$-means [26], with an iterative algorithm converging in $I$ steps. We note that this method is not memory conservative, as we need to store all data-points to recompute the cluster information. However, with warm-starts, it can be quite computationally efficient; while the theoretical upper bound on the time-complexity is unchanged compared to not warm-starting, in practice, warm-starting greatly reduces the required number of iterations for convergence.
In the complexity analysis, we assume the $k$-centers approximation to be $k$-means [26], with an iterative algorithm converging in $I$ steps. This method is not memory conservative, as we need to store all data-points to recompute the cluster information. However, with warm-starts, it can be quite computationally efficient; in practice, warm-starting greatly reduces the required number of iterations.
@@ 6.3 Online clustering (OnlineClustering)
For an algorithm that is also memory conservative, we propose an online updating procedure that takes inspiration from a well-known online clustering approach, CluStream [2]. We summarize the algorithm here, and give a more detailed version in Appendix A.4. Similarly to CluStream, we maintain two sets of clusters: a set of $Q\geq K$ microclusters, and a set of $K$ macroclusters. Data-points are assigned to microclusters on arrival, or initialized as a new cluster based on its distance to current clusters. When new microclusters are initialized, the closest microclusters are merged to maintain the budget $Q$. The $K$ macroclusters are obtained by clustering the microclusters, and can be warm-started.
Differently from CluStream, we allow for separate cluster centers and centroids; the centers do not shift with the addition of new data-points, while the centroids are updated. For times $t\leq\tau$, all clustering assignments can be made according to either the centers or the centroids. After $t>\tau$, cluster assignments must be made according to the centers of the $K$ macroclusters. We require the macrocluster update to also be performed online, so that they have up-to-date information. For all clusters, we formalize the notion of the cluster supports, such that they are consistent with Definition 3.1 and Assumption 3.2. This requires adjustments to the Voronoi regions around the centers/centroids, based on the already-clustered data-points. For continuous distributions ${\mathbf{P}}$, since the data-points have measure 0, this does not need to be explicitly maintained. Therefore, this algorithm can discard all data-points once seen, and keep only the aggregated information of size $O((K+Q)d)$. For discrete ${\mathbf{P}}$, however, the adjustment to the support is no longer negligible. Nevertheless, discrete distributions generally have a finite set of atoms, so memory is not a bottleneck.
For the complexity analysis, we assume the $k$-centers approximation on the $Q$ microclusters to be $k$-means [26]. Compared to the reclustering approach, we are solving a smaller problem, which reduces the theoretical time complexity; the set of $Q$ microclusters allows this method to interpolate between SupCover and Reclustering. Practically, however, the main advantage of this approach is the lower memory requirement.
@@ 7 Choosing hyper-parameter values
Above, we detailed various theoretical bounds for the online algorithm, with a given $K$ and with $\varepsilon^{K}_{t}$ chosen according to measure concentration results. In practice, however, these values are hyper-parameters that can be tuned. In this section, we give some guidelines for their practical selection.
@@
As noted in the previous sections, the value of $K$ controls the discrepancy between the online and nominal DRO performance. A low value of $K$ is computationally efficient, while higher values allow for better approximations of the underlying distribution, thus increasing solution quality. In practice, we expect a value of $K$ that doesn’t far exceed the number of modes of the underlying distribution. If the initial dataset $\mathcal{D}_{0}$ is large enough, this value of $K$ can be chosen a-priori, using the elbow method on the clustering value $(D^{K}_{0,2})^{2}$ [49]. If the initial dataset is small, we can initialize with a small value of $K$, and allow it to increase or decrease, based on its performance. Our asymptotic analysis requires only that after a finite time $\tau$, the cluster supports are fixed; prior to this time, the clusters are flexible, and could be adjusted to better approximate the true distribution.
@@
In Wasserstein DRO, for a tighter practical certificate on the out-of-sample performance, the radius $\varepsilon$ is often chosen through cross-validation. We employ a similar technique for the compressed online algorithm, where we tune $\varepsilon^{K}_{t}$ to minimize the out-of-sample performance. We assume to have a validation dataset of size $N^{\rm val}$, in addition to our growing training data. For every time step $t$, upon constructing the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$, we solve the optimization problem (8) with a finite set of radii $\varepsilon^{K}_{t}$. For each in-sample solution, we use the validation dataset to estimate the out-of-sample performance via SAA, and record whether or not it is upper bounded by the certificate $H^{K}_{t}+\underline{\psi}_{t}^{K}$, defined in Theorem 3.5. We select the sequence of $\varepsilon^{K}_{t}$ that gives the lowest out-of-sample performance, and report the corresponding sequence of data-driven solutions, certificates, and empirical confidence.

### v2

@@ 4 Special case: bounded support
In the dimension-dependent portion of Section 2.2, we gave performance guarantees for light-tailed supports $S$, and set the ambiguity set radius $\varepsilon^{K}_{t}$ with the nominal DRO radius $\varepsilon_{t}$, calculated using Theorem 2.1. In this section, for the case of bounded support, we give an explicit formulation of the radius in terms of the diameter of the support $S$. We also derive performance guarantees using a different theoretical approach from that of Section 3.3: using measure concentration results, we instead increase the ambiguity set radius by a finite value. In this case, the online objective value $H^{K}_{t}$ is guaranteed to be an upper bound on the DRO objective value $H_{t}$, regardless of the curvature of $f$.
@@ 4.1 Wasserstein radius with bounded support
This follows from [10, Proposition A.2] and exploits the fact that bounded distributions are sub-Gaussians. Then, by bounding the expected Wasserstein distance using a function of $d,p,\tilde{\rho}$, and $N$, we obtain the following explicit characterization of the nominal ambiguity radius. For details, see the proofs in the online version of the referenced paper.
@@
For any time step $t$ and $\beta_{t}$, the radius $\varepsilon_{t}$ calculated from Theorem 4.1 can be used to obtain the performance guarantees in Theorem 2.1, since distributions with bounded support are trivially light-tailed.
@@ 4.2 Performance guarantees with bounded support
Let $\hat{{\mathbf{P}}}^{K}_{t}$ be the clustered empirical distribution at time $t$, with cluster diameters $\mathrm{diam}_{q}(C^{k}_{t})\leq 2\eta_{K}$, defined according to Definition 3.5 and an arbitrary norm $q$. Then, for all $t$ and $p$, and with the same norm, we have $W^{K}_{t,p}\leq 2\eta_{K}.$ With probability $1-\beta_{t}$, we have $W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})\leq\varepsilon_{t}+2\eta_{K},$ where $\beta_{t}$ and $\varepsilon_{t}$ are computed as in Theorem 4.1.
@@
Under the assumptions of Theorem 4.1, let the adaptive ambiguity set $\mathcal{P}^{K}_{t}$ be defined with $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$. Then, the solution $x^{K}_{t}$ and optimal value $H^{K}_{t}$ of the compressed DRO problem (7) implies the finite sample performance guarantee
${\mathbf{P}}^{t}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x^{K}_{t})]\leq H^{K}_{t}\right)\geq 1-\beta_{t}.$
@@
This follows from Lemma 4.1 and the same logic as nominal DRO results [38, Theorem 3.5]. ∎
In addition, we can derive the following asymptotic result by applying a similar approach as [38, Lemma 3.7].
@@
This lemma dictates that asymptotically, the true distribution lies in a ball around the clustered distribution with radius $4\eta_{K}$; the increased radius is the price to pay due to the limited computational budget available to solve (4), and is directly related to the quality of our approximation $\hat{{\mathbf{P}}}^{K}_{t}$. Clearly, if $K$ is high, then $\eta_{K}$ will be low and our asymptotic estimate will be more accurate; viceversa for low $K$. Following [38, Theorem 3.6], we now derive the following result.
@@
Assume for all $t$, $\varepsilon_{t}$ is computed as in Theorem 4.1, with a sequence $\beta_{t}$ such that $\sum_{t=0}^{\infty}\beta_{t}<\infty$ and $\lim_{t\to\infty}\varepsilon_{t}=0$. Also assume the adaptive ambiguity set $\mathcal{P}^{K}_{t}$ is defined with $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$, and that for all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ satisfy the $M$-Lipschitz condition given in Assumption 2.1 with constants $M_{j}$. Then, the sequence of solutions and optimal values $x^{K}_{t}$ and $H^{K}_{t}$ of the compressed DRO problem (7) satisfies
@@ 5 Online algorithms for data-compression
In Section 3.1, we gave requirements for the clustering algorithm. In this section, we provide three online clustering algorithms satisfying Assumption 3.1. These algorithms maintain a reasonably accurate approximation of the optimal discretization while minimizing computational effort. Note that they can be easily adjusted to perform updates with a batch of data, instead of a single datapoint, in case the data-arrival frequency is high. Below, we summarize the time and memory complexities of the algorithms.
Table 1: Time and memory complexities of our clustering algorithms. Recall that $n_{t}$ is the total number of datapoints at time $t$, $K$ is the number of clusters, and $d$ is the dimension of the data. We also let II be the number of iterations required for the clustering algorithm’s convergence. The values $G$ and $Q$ are defined in the respective algorithms.
@@ 5.1 A bounded coverage of the support (SupCover)
3:assign datapoints to clusters $\{C^{k}_{0}\}_{k=1}^{K}$, and compute $\{n^{k}_{0}\}_{k=1}^{K}$, $\{\theta^{k}_{0}\}_{k=1}^{K}$, $\{\bar{u}_{0}^{k}\}_{k=1}^{K}$
@@
6:  $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\} \triangleright$ assign $\hat{u}$ to cluster $k^{\prime}$ following  (18)
@@
8:  $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$ $\triangleright$ update mean
@@
where $\hat{u}$ is the datapoint at time $t$. Additionally at each time $t$, the mean of the cluster that gained a datapoint is updated. Notably, we do not need to store $\mathcal{D}_{t}$. By keeping track of the centers, means and weights of each cluster, together with a count of the total number of datapoints seen, we can perform the above updates with only the newest datapoint. After performing the update, this datapoint can also be discarded. The algorithm is thus memory efficient; it only requires $O(Kd)$ space.
@@ 5.2 Approximate $k$-centers clustering with warm starts (Reclustering)
[…]
3:assign datapoints to clusters $\{C^{k}_{0}\}_{k=1}^{K}$, and compute $\{n^{k}_{0}\}_{k=1}^{K}$, $\{\theta^{k}_{0}\}_{k=1}^{K}$, $\{\bar{u}_{0}^{k}\}_{k=1}^{K}$
@@
8:   $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\} \triangleright$ assign $\hat{u}$ to cluster $k^{\prime}$ following  (18)
9:   $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$ $\triangleright$ update mean
@@
In the complexity analysis in Table 1, we assume the $k$-centers approximation to be $k$-means [23], with an iterative algorithm converging in $I$ steps. This method is not memory conservative, as we need to store all datapoints to recompute the cluster information. However, with warm-starts, it can be quite computationally efficient; in practice, warm-starting greatly reduces the required number of iterations.
@@ 5.3 Online clustering (OnlineClustering)
For an algorithm that is also memory conservative, we propose an online updating procedure that takes inspiration from a well-known online clustering approach, CluStream [2]. We summarize the algorithm here, and give a more detailed version in Appendix A.3. Similarly to CluStream, we maintain two sets of clusters: a set of $Q\geq K$ microclusters, and a set of $K$ macroclusters. Datapoints are assigned to microclusters on arrival, or initialized as a new cluster based on its distance to current clusters. When new microclusters are initialized, the closest microclusters are merged to maintain the budget $Q$. The $K$ macroclusters are obtained by clustering the microclusters, and can be warm-started.
Differently from CluStream, we allow for separate cluster centers and means; the centers do not shift with the addition of new datapoints, while the means are updated. Both values are updated when clusters are merged. For times $t\leq\tau$, all clustering assignments can be made according to either the microcluster centers or means. We also require the subsequent macrocluster update to be performed online, either with the addition of each datapoint, or before each time the DRO problem (9) is solved. For $t>\tau$, cluster assignments must be made according to the macrocluster centers, which are fixed at time $\tau$. For all clusters, the support is the Voronoi region around its center/mean, with adjustments based on the already-clustered datapoints, which may fall outside the region. For continuous distributions $\mathbf{P}$}, since the datapoints have measure 0, this does not need to be explicitly maintained. Therefore, this algorithm can discard all datapoints once seen, and keep only the aggregated information of size $O((K+Q)d)$. For discrete $\mathbf{P}$}, however, the adjustment to the support is no longer negligible. Nevertheless, discrete distributions generally have a finite set of atoms, so memory (for storing the support information and datapoints) is not a bottleneck.
For the complexity analysis, we again assume the $k$-centers approximation on the $Q$ microclusters to be $k$-means [23]. Compared to the reclustering approach, we are solving a smaller problem, which reduces the theoretical time complexity; the set of $Q$ microclusters allows this method to interpolate between SupCover and Reclustering. Practically, however, the main advantage of this approach is the lower memory requirement.
@@ 6 Choosing hyper-parameter values
Above, we detailed various theoretical bounds for the online algorithm, with a given $K$ and with $\varepsilon^{K}_{t}$ chosen according to measure concentration results. In practice, however, we tune these hyper-parameters based on the performance on a validation set of size $N^{\rm val}$.
@@
If the initial dataset $\mathcal{D}_{0}$ is large enough, the value of $K$ can be chosen a-priori, using the elbow method on the clustering value $(D^{K}_{0,2})^{2}$ [46, 54]. If the initial dataset is small, we can initialize with a small value of $K$ and adapt it based on the performance on validation data.
@@
Similarly to existing literature on Wasserstein DRO, we tune radius $\varepsilon_{t}$ to minimize validation performance: for every time step $t$, upon constructing the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$, we solve the optimization problem (9) with a finite set of candidate radii $\varepsilon^{K}_{t}$. For each solution, we use the validation dataset to compute the validation performance. Averaging across multiple repetitions of the experiment, we select the sequence of $\varepsilon^{K}_{t}$ that gives the best validation performance.